The function g left(xright)=fracx−−3 is not defined for x=___________________.
a.7
b.\\(frac{5}{7}\\)
c.5
d.\\(frac{7}{5}\\)
1
Expert's answer
2020-05-22T16:16:28-0400

Note: Given function is "g(x)=\frac{x−3}{5x-7}" .

Given function is not defined when the denumerator is zero.

Denumerator is zero when "5x-7=0" "\implies 5x = 7 \implies x =\frac{7}{5}".

Hence the given function is not defined at "x = \frac{7}{5}".

Correct option is option (d).